Use the Laplace transform to solve the given initial-value problem. y'' + y = sqrt2 sin( sqrt2 (t)), y(0) = 5, y'(0) = 0 y(t) =
Answer #1

y+y = 5 Sim (15+) By arelying balloce Transform Ś y cs) -SYC) –Y 6) + YCS) 5. C) Griver, XCO) = 5 y(o) Yes) Cs?+1) - 55 =